OpenStudy (anonymous):

Solve the system using substitution. 3x + 2y = 23 1/2x - 4 = y

OpenStudy (anonymous):

replace \(y\) in the first equation by \(\frac{1}{2}x-4\) ( because that is what it is) and solve \[3x+2\left(\frac{1}{2}x-4\right)=23\] for \(x\)

OpenStudy (mathmale):

This is solution by substitution. substitution has eliminated the variable y, making it that much easier to solve for x.

OpenStudy (anonymous):

\[3x+2\left(\frac{1}{2}x-4\right)=23\]\[3x+x-8=23\] is a start

OpenStudy (triciaal):

@KaaayKillaaah please review your response. What are you thinking?

OpenStudy (triciaal):

@trentren follow what satelite73 has

OpenStudy (anonymous):

Okay I got 4x = 31 @satellite73

OpenStudy (anonymous):

looks right to me

OpenStudy (anonymous):

Its not an answer choice though.

OpenStudy (anonymous):

that is not an answer, that is a step along the way

OpenStudy (anonymous):

if \[4x=31\]what is \(x\)?

OpenStudy (anonymous):

x = 7.75

OpenStudy (kaaaykillaaah):

@triciaal .. I know what I am doing @satellite73 did the same thing

OpenStudy (anonymous):

ok now find \(y\) if the answer choices have all whole numbers, then perhaps there is a typo in the question if one of the answer choices has \(x=7.75\) then we are in good shape

OpenStudy (anonymous):

These are the answer choices a. ( 9, -2 ) b. ( 27/4, -5/8 ) c. ( 31/4, -1/8 ) d. ( 3/2, -13/4 )

OpenStudy (anonymous):

16?

OpenStudy (triciaal):